Hi what is 5x + 3 = 7x -1
Isolate x

Found 2 solutions by ikleyn, math_helper:
Answer by ikleyn(53763)

.

To isolate x, move term 5x from the left side to the right side, changing the sign to opposite one;


at the same time move the term (-1) from the right side to the left side, changing its sign to the opposite one.


Then you get this equivalent equation


    3 + 1 = 7x - 5x


and "isolation" is completed.

Solved.

-------------

If you struggle solving linear equations, look into the lesson
    - HOW TO solve a linear equation
in this site, and learn the subject from there.


Happy learning (!)



Answer by math_helper(2461)


5x + 3 = 7x - 1
I will show every step, but with practice, you will "move" terms from
one side to the other.
Subtract 5x from each side:
5x + 3 - 5x = 7x - 1 - 5x
Combine like terms (on the left 5x-5x = 0x which we just drop completely, on the right 7x-5x = (7-5)x = 2x):
3 = 2x - 1
Add 1 to each side:
3 + 1 = 2x - 1 + 1

Combine:
4 = 2x

Divide both sides by 2:
4/2 = 2x/2 (2x/2 = 1x which we write simply as 'x', next line)
Simplify:
2 = x ---> Ans: x = 2

----------
With experience, you would either do the whole thing in your head or you might do one or two steps:
5x + 3 = 7x - 1
4 = 2x (added 1 to both sides, also subtracted 5x from both sides)
2 = x (divided both sides by 2)

Solving for x/isolating x is based on this simple principle: if there is an equation, then doing the same valid operation to both sides does not affect that equality.